1 2 1 Show that A=0 0 a diagonal D. 2 is diagonalizable by finding P and D such that P-1AP = D for 3 0
Answer #1

Solution: Given

A= 1 1 0 0 2 2 0 0 3

Now eigen values are given by

A - XI = 0

\Rightarrow \begin{vmatrix} 1-\lambda & 1 & 0\\ 0 & 2-\lambda & 2\\ 0 & 0 &3-\lambda \end{vmatrix}=0

\Rightarrow(1-\lambda)(2-\lambda)(3-\lambda )=0

\Rightarrow\lambda=1,2,3

all eigen values are distinct and real so A is diagonalizable.

Now, invertible matrix P is given by the eigenvectors.

So let us find eigen vectors corresponding to each eigen value.

Now, eigen vector v_1 corresponsing to \lambda=1 is given by

[A-I]v_1=0

\Rightarrow \begin{bmatrix} 0 & 1 &0 \\ 0 & 1 &2 \\ 0 &0 & 2 \end{bmatrix}\begin{bmatrix} x\\y \\ z \end{bmatrix}=\begin{bmatrix} 0\\0 \\0 \end{bmatrix}

> y=0,2 = 0

By choosing x=1 , we get

v_1=\begin{bmatrix} 1\\0 \\ 0 \end{bmatrix}

Now, eigen vector v_2 corresponsing to \lambda=2 is given by

[A-2I]v_2=0

\Rightarrow \begin{bmatrix} -1 & 1 &0 \\ 0 & 0 &2 \\ 0 &0 & 1 \end{bmatrix}\begin{bmatrix} x\\y \\ z \end{bmatrix}=\begin{bmatrix} 0\\0 \\0 \end{bmatrix}

\Rightarrow x=y,z=0

By choosing x=1 , we get

v_2=\begin{bmatrix} 1\\1 \\ 0 \end{bmatrix}

Now, eigen vector v_3 corresponsing to \lambda=3 is given by

[A-3I]v_3=0

\Rightarrow \begin{bmatrix} -2 & 1 &0 \\ 0 & -1 &2 \\ 0 &0 & 0 \end{bmatrix}\begin{bmatrix} x\\y \\ z \end{bmatrix}=\begin{bmatrix} 0\\0 \\0 \end{bmatrix}

\Rightarrow 2x-y=0,\ \ \ y-2z=0

\Rightarrow 2x=y,\ \ \ y=2z

\Rightarrow x=z,y=2z

By choosing z=1 , we get

v_3=\begin{bmatrix} 1\\2 \\ 1 \end{bmatrix}

Thus,

P=\begin{bmatrix} 1 & 1&1 \\ 0& 1& 2\\ 0& 0 & 1 \end{bmatrix}

Now find inverse of P .

Let

P=IP

\Rightarrow \begin{bmatrix} 1 & 1&1 \\ 0& 1& 2\\ 0& 0 & 1 \end{bmatrix}= \begin{bmatrix} 1 & 0&0 \\ 0& 1& 0\\ 0& 0 & 1 \end{bmatrix}P

applying R_1\to R_1-R_3,\ R_2\to R_2-2R_1

\Rightarrow \begin{bmatrix} 1 & 1&0 \\ 0& 1& 0\\ 0& 0 & 1 \end{bmatrix}= \begin{bmatrix} 1 & 0&-1 \\ 0& 1& -2\\ 0& 0 & 1 \end{bmatrix}P

applying R_1\to R_1-R_2

\Rightarrow \begin{bmatrix} 1 & 0&0 \\ 0& 1& 0\\ 0& 0 & 1 \end{bmatrix}= \begin{bmatrix} 1 & -1&1 \\ 0& 1& -2\\ 0& 0 & 1 \end{bmatrix}P

compare it with

I=P^{-1}P

we get

P^{-1}= \begin{bmatrix} 1 & -1&1 \\ 0& 1& -2\\ 0& 0 & 1 \end{bmatrix}

Hence,

D=P^{-1}AP

\Rightarrow D=\begin{bmatrix} 1 & -1&1 \\ 0& 1& -2\\ 0& 0 & 1 \end{bmatrix} \begin{bmatrix} 1 & 1 & 0\\ 0 & 2 & 2\\ 0 & 0 &3 \end{bmatrix}\begin{bmatrix} 1 & 1&1 \\ 0& 1& 2\\ 0& 0 & 1 \end{bmatrix}

\Rightarrow D=\begin{bmatrix} 1 & -1&1 \\ 0& 1& -2\\ 0& 0 & 1 \end{bmatrix} \begin{bmatrix} 1 & 2 & 3\\ 0 & 2 & 6\\ 0 & 0 &3 \end{bmatrix}

\Rightarrow D=\begin{bmatrix} 1 & 0&0 \\ 0& 2& 0\\ 0& 0 & 3 \end{bmatrix}

Which is the required solution.

This complete the solution.